### v1

Abstract
TikTok is now a massive platform, and has a deep impact on global events. But for all the preliminary studies done on it, there are still issues with determining fundamental characteristics of the platform. We develop a method to extract a representative sample from a specific time range on TikTok, and use it to collect >99% of posts from a full hour on the platform, alongside a dataset of >99% of posts from a single minute from each hour of a day. Through this, we obtain post metadata, video media data, and comments from a close to complete slice of TikTok. Using this dataset, we report the critical statistics of the platform, notably estimating a total of 117 million posts produced on the day we looked at on TikTok.
keywords:
tiktok, sample, social-media, vertical-slice
1 Introduction
TikTok has reshaped global culture, marketing, and entertainment through its highly engaging, algorithm-driven platform that promotes short-form video content. Due to its immense reach [1], especially among younger voters [2, 3], and its ability to amplify content quickly through its algorithm, politicians, advocacy groups, and individuals use TikTok to share political messages, mobilize voters, and raise awareness about issues [4, 5].
However, concerns have been raised about the spread of misinformation [6], the platform’s potential for algorithmic bias [7], the lack of transparency in how political content is managed [8], and child safety on the platform [9]. Additionally, TikTok’s ownership and data practices have sparked debates over national security and foreign influence, particularly regarding its potential use in shaping public opinion during elections [10, 11, 12]. At the same time, little is known about the platform’s inner working and its global activity level. If we want to gain more insight into how TikTok feeds into these issues, we need strong quantitative insight into the platform.
Data collection methods on TikTok are limited. Search mechanisms on the site are limited to hashtag based search which stops at 1000 returns, fuzzy search, and user and song based search [13]. All of these methods limit the extent to which a true complete slice of the platform can be examined. While TikTok has been offering a Researcher API to selected scholars [14], it has been shown recently, that this API shows significant deviations in terms of metadata and availability of videos when compared to the TikTok website [15]. Additional limitations of the API include restrictive user agreements [16], and the inability to return videos originating from Canada  [^1] https://developers.tiktok.com/doc/research-api-codebook/. We must obtain stronger sampling methods if we are to gain better insight into a platform of TikTok’s magnitude [17]. Without a representative sample from such a method, researchers cannot properly contextualize or validate findings about behaviour on TikTok.
To address this need, we developed a novel method for collecting >99% uploaded to TikTok within a specific timespan, enabling a collection of a close to complete slice of TikTok for the first time.
We did this by reverse engineering how TikTok creates unique identifiers (IDs) for its videos, in order to create an ID generator that can allow us to search for all possible TikTok videos, which allows us to bypass the black box search elements normally available. Previous work using this method has allowed a greater understanding of YouTube [18].
Considering the focus of research on TikTok and social media platforms in general, we will provide insight on the following points:
  • The total volume of content on TikTok: Estimating this allows us to understand the true size of the platform for the first time, and therefore understand its total usage, a critical task to avoid sampling bias [19].
  • Distributions of post content on TikTok: We present data on country post origin, engagement statistics, and local posting time from our representative dataset. This allows us to understand global usage and consumption patterns, critical when considering the disparity between global populations and the TikTok user base [20].
  • Media content: Video media content is the central medium of TikTok, and understanding video topic breakdowns is therefore critical for insight into what role topics have on the platform. Additionally, TikTok has proven incredibly popular for children [21], raising safety concerns, so understanding the role of children in content on the platform is essential.
  • Comment statistics: Comment volume, comment posting time, and comment language. TikTok comments give us insight into mass user engagement patterns.
As part of this work, we will release the fundamental statistics necessary to understand TikTok, and calibrate studies done on it. This data is critical for grounding findings from TikTok. With that in mind, we will find these statistics with a methodology detailed in this work, and release them at URL.
With this study, we have violated TikTok’s Terms of Service  [^2] https://www.tiktok.com/legal/page/us/terms-of-service/en, specifically the following section: ‘use automated scripts to collect information from or otherwise interact with the Services;’ However, we strongly believe that our work is critical to better understand a platform that is used by billions of users worldwide. Greater platform transparency will lead to higher quality and more trustworthy research, which in turn, will help the community and TikTok work together to help the platform flourish.
2 Results
We developed a method that, to the best of our knowledge, allows us to collect >99% of the post content on TikTok across the world. We used that method to collect every video made between 5-6pm UTC on the 10th April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all videos from the 42nd minute of each hour from 10pm UTC on the 9th April, till 10pm on the 10th April, to obtain across day representative data. Here, we use that data to extract what we consider to be the critical statistics of note on TikTok, that will let the research community contextualize findings they have made on TikTok.
2.1 Statistics
We first present the high level statistics of our two datasets, enabling comparison with other dataset sizes.
Metric 1 Hour 1 Minute Each Hour
Number of Videos 5,157,488 1,879,335
Number of Unique Users 3,715,672 1,760,335
Average Comment Count 4.92 4.18
Average Like Count 177.9 145.8
Average Share Count 14.1 9.14
Average View Count 2,533.2 2243.9
Average Video Duration (seconds) 20.69 20.42
Table 1: Statistics from the two datasets we collect for this work. Note that the two datasets overlap at the $42^{nd}$ minute of the hour we looked at in this work, so total counts are lower than the sum/mean of the two counts.
(a) Histogram of comment count.
(b) Histogram of view count.
(c) Histogram of like count.
(d) Histogram of share count.
(e) Histogram of video duration.
Figure 1: Histograms of various statistics in the video metadata, across the 1 hour dataset.
Figure 1 shows the distribution of video statistics from the 1 hour dataset we collected. While comment count and share count follow a typical exponential distribution as previously seen [14], view count and like count show a mode at $\sim$100 views and $\sim$50 likes respectively, differing from the distributions reported by Corso et al. [14]. This may be due to the large discrepancies between the Research API and the data returned to the platform interface demonstrated by Pearson et al. [15], highlighting the need to double check Research API data with main platform data.
2.1.1 Estimating overall post volume
The nature of our dataset gives us first insight into the overall size of TikTok, a critical piece of knowledge for any social media platform [17].
For the observed hour on TikTok, we collected $5,157,488$ videos with an average of $1,432.6$ videos [$1,166-1,758$] per second (see Fig. 2(a)). The aggregated numbers per minute range from $80,292$ to $95,915$ (avg. $85,958.1$) videos per minute (Fig. 2(b)).
In order to estimate the videos per hour for an entire day, we collected the videos of the $42^{nd}$ minute of every hour. For our collected hour, the $42^{nd}$ minute is slightly less active ($\sim 96.5\%$) than the average of the other $59$ minutes. Consequently, we used this correction factor to calculate the estimated number of videos per hour. The result can be seen in Fig. 2(c) and shows an interesting two-peaked temporal pattern. These estimated hours aggregate to 116.7 million video posts over a 24-hour time period.
(a) Posts per second.
(b) Posts per minute.
(c) Estimated posts per hour.
Figure 2: TikTok posts per second and per minute on the 10th of April, 2024, 5pm–6pm UTC, as well as the estimated posts per hour over a 24 hour time period.
We note that the time distribution of videos per hour looks markedly different from Corso et al. [14]. We also see that there is a 14% higher posting volume at the 0th second of each minute (where the distribution is otherwise flat), implying bot activity, or scheduled posts. View counts for videos posted at the zeroth second get significantly more views (t-test 3.97, p-value: 7e-5), with an average of 3899 views at the zeroth second vs. 2502 views for all other seconds. Videos at the zeroth second more frequently use hashtags implying the desire to reach a broader audience, such as fyp, implying that these posters are using post scheduling and hashtag use to attempt to reach a broader audience.
Understanding the total size of the platform gives much needed context when we want to understand samples of it.
2.1.2 Understanding global usage
Figure 3: Choropleth of video posting per capita for all countries, determined via the locationCreated tag available on the video post metadata. We normalize by the population of each country to show posts per capita. The counts are corrected for geographical regions that we estimate that we underrepresent.
Figure 4: Regression of country population against video count in this dataset, with the 6 countries furthest from the trend line labelled, as determined by highest and lowest residuals from the linear regression.
We then looked at global usage of TikTok. Research on TikTok frequently looks at country specific slices [22], so understanding actual country level usage is critical for context [20]. We infer post country creation from the locationCreated tag that refers to video origin country  [^3] https://developers.tiktok.com/doc/research-api-specs-query-videos. While this indicator of country origin is likely confounded by VPN usage and other factors, it represents our current best method for assessing post country origin. Our method has slight variation in regional representativeness, so we adjust counts based on known statistics from our method, detailed in Section 4.3.
Figure 3 shows the popularity of TikTok across the world, correcting data by the population of the country  [^4] https://data.worldbank.org/indicator/SP.POP.TOTL.
Figure 4 shows a regression plot of country population against video count in our dataset. We see a strong fit to the data, evidence that we have successfully collected a globally representative dataset. We see that China, India, and Cuba, have the least videos in our dataset per capita, based on the residual of a fitted linear regression. We note that while TikTok was previously extremely popular in India [14], there are now very few videos originating from India (as deemed by TikTok), likely due to the nationwide censorship policy on Chinese applications there [23]. Of course, there could still be videos produced from India via virtual private networks (VPNs), which would be tagged with a different origin country by TikTok. TikTok is not officially available in China [24], and Cuba has heavy internet restrictions [25]. Monaco, Qatar, and the UAE, have the most videos in our dataset per capita.
Using the locationCreated attribute, we can also infer the distribution of when TikTok posts are created in the local time, see Figure 5. In the case of countries spanning multiple timezones, we simply select the first timezone associated with a country from the Olson timezone database  [^5] https://github.com/stub42/pytz/blob/master/tz/zone.tab. Inaccuracies are inevitable due to lack of sub-country location information, and countries spanning multiple timezones, but the general trends should be preserved nonetheless.
Figure 5: Distribution of local time video posting frequency, found by looking at video post time combined with the locationCreated tag.
TikTok’s global popularity means that understanding TikTok’s relative popularity across the world is necessary when trying to understand local aspects of the system, so we release country statistic breakdowns in our data release.
2.2 Video Media
TikTok is a video based platform, and any understanding of the nature of the platform as a whole needs to come with a look at the actual video media data of the platform.
For each post in our dataset that has associated video media data (i.e. not a story-post), we downloaded the video media data. We wanted to gain a view into what kind of topics were discussed on the platform. We therefore used a combination of topic modelling and vision language models to cluster the videos, and describe each cluster. We plot the posts and clusters in Figure 6.
Figure 6: View of the topics covered by the video posts, with each small dot representing a video, and distance between dots representing video content similarity. Common topics of videos are colored and labelled with a description describing the topic.
We can see the broad array of topics with the largest being videos of a person talking directly to the camera, along with common activities in everyday life such as sports, transport, and lifestyle videos. In smaller clusters we find more uncommon topics, such as ‘Political Commentary on Police Operations in Pakistan’ (0.025% of videos), ‘Camouflaged Military Soldier Image’ (0.06%), and ‘Palestinian Flag Protest Imagery’ (0.008%).
Now, when we look at specific activity on TikTok, such as political, military, and healthcare content, we can start to understand what proportion of the whole platform it represents.
2.2.1 Presence of Children
The heavy use of TikTok by children and teenagers has caused concern for safety on the platform. But the lack of a representative sample of TikTok has meant so far we haven’t been able to estimate the true magnitude of child activity on the platform. We know there is a high prevalence of children on the platform, with 50% of 5-7 year olds, and 80% of 12-15 year olds in the UK using the platform [21], but we don’t what percentage of content on the platform features children. We aim to address that here.
We trained a classifier to detect if a video contains a child, with a mean F1 of 0.86 from 5-fold cross-validation on our train/testset (precision 0.88, recall 0.81). We used this classifier to find that 18.13% of the videos in our 24 hour dataset contain children (95% confidence interval: [18.06%, 18.21%] from bootstrapped confidence intervals). This is a strikingly high statistic, and we believe there is not a specific topic for children in our topic model purely because there they appear so commonly and diversely in the videos. We then wanted to know how this measure breaks down by country.
We removed countries where there were insufficient post counts to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05).
Figure 7: Choropleth of share of videos from our 1 minute per hour dataset containing one or more children, divided by the share of the population under the age of 14 [26] (to avoid showing simply a map of countries with a youthful population), plotted by country. Solid areas are statistically significant, hatched areas are not significant as determined by a binomial test (p=0.05).
Figure 8: Scatter plot of countries, for each country showing the share of the population under the age of 14 against the share of videos from that country containing one or more children, as determined by our video classifier. We have removed countries where there was insufficient data to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05). We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively.
Figure 7 shows a choropleth of share of videos containing one or more children per country, adjusted by the share of the population under the age of 14  [^6] https://population.un.org/wpp/ (to avoid showing simply a map of countries with a youthful population). This effectively shows countries where there is a culture of posting videos containing children. Figure 8 shows a scatter plot of countries comparing the share of the population under 14, and the share of videos containing one or more children, including a linear regression. We label the top 3 and bottom 3 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively. This data therefore lets us couch our understanding of the prominence of children on TikTok in the actual fraction of the platform they take up.
With this analysis, we produced, to the best of our knowledge, the first attempt to measure what percentage of TikTok consists of content featuring children, lending context to our understanding of the role children play on TikTok.
2.3 Comments
Comments are the richest form of mass engagement data on TikTok, providing language, semantic, and temporal data. We can use them to build understanding of TikTok’s audience.
We collected 18.4 million comments directed at the videos (excluding comments on comments), associated with 13.5 million commenters. We collected these comments in September 2024, 5 months after the day the videos were posted, using Ensemble Data  [^7] https://ensembledata.com/.
The median number of comments on the videos is 0, and the mean number is 4.92, with 51.9% of videos receiving zero comments.
When looking at the time distribution of comment creation time, we see a flat distribution (std dev. across all seconds: 652, std dev. across all minutes: 3.7k, with a mean of 306k), indicating minimal scheduled posting of comments, and if automated comment posting is occurring, timing of posting is more sophisticated.
We also looked at average offset between video post and comment time, as shown in Figure 9. We see that for all commenters, the modal time to comment is just under 1 day, and that more frequent commenters comment have a modal peak of 1 hour from post time. We can also see that around 2/3 months after a video is posted, the commenting frequency on the video rapidly declines. This data shows us that while distribution processes on TikTok are fast, users are not engaging with the content until at least a few minutes past post time.
Looking at comment language, 47% of comments are unknown (likely due to the high prevalence of emoji only comments on the platform), 15% are in English, 9% are in Spanish, and 5% are in Arabic. The high percentage of emoji only commenting on TikTok could imply large quantities of inauthentic activity, as emojis would be an easy way to boost a video’s engagement while requiring limited effort on behalf of the bot.
Figure 9: Time offset between video and comment creation by commenter frequency. We show offset distributions for authors who have at last 1 comment, authors who have at least 5 comments, and authors who have at least 25 comments, to show how commenting frequency affects commenting speed.
We will release the distributions from these analyses in our data release to allow other members of the community to use them to add context to their own work.
2.4 Error Analysis
A key feature of our collection method means that we receive a great number of error messages when we attempt to collect posts that either don’t exist, or are not available for other reasons. As to our knowledge this has not previously been reported on, we report on it here.
We received several different types of error message for a requested video, and we report the most common error types, excluding the error message ‘item doesn’t exist’, in Table 2. We can see that an extremely large number of video posts are deleted, or privatized (status_self_see). We also gain an insight into what seems to be the TikTok moderation process, with  7% of videos unavailable at the time of collection due to a review stage (status_reviewing),  1% of videos hidden after audit (status_audit_not_pass), and a small number potentially currently undergoing content classification for moderation (content_classification). We also see two error messages potentially related to geo-location, with cross_border_violation likely related to location based content restrictions, one of which was deployed around the time of the invasion of Ukraine [27].
Status Count Pct. (%)
status_deleted 3,763,773 20.96
status_self_see 3,732,155 20.79
status_reviewing 1,201,251 6.69
status_audit_not_pass 191,690 1.07
content_classification 26,654 0.15
cross_border_violation 2,134 0.01
copyright_geo_filter 296 0.00
Table 2: Distribution of most common post error messages, excluding ‘item doesn’t exist’. Reported percentage is out of all videos that at some point existed in our combined datasets. So from this we can infer that around 20% of videos on TikTok are deleted.
When we look at deletion rates over different parts of our collection period, we see that of the posts that exist, at 26 days after post time, deleted videos accounted for 12% of videos, and at 215 days after post time, they accounted for 20% of videos.
This data gives us a first insight into inaccessible content on TikTok, and future work can exploit this to bring to light the moderation processes used by the platform, the rate of deletion, and the nature of cross-border content restrictions.
2.5 Inauthentic Activity
With any social media dataset, understanding human behaviour through it also requires understanding how much of the data is produced through inauthentic activity [28]. Given that there are no publicly available bot detection tools for TikTok, we used simple heuristics for bot detection [29]. We find that there are 46 accounts with more than 100 videos each posted in the hour we looked at, and the top 5 highest frequency posting accounts, with between 1500 and 6000 videos in the hour we looked at, all have the default username of user followed by 9 random digits. Their videos have an average of 1.9 views each, but account for 0.28% of the total 1 hour dataset. Looking at comments, there are 831 accounts that have at least 10 comments and whose mean commenting frequency is higher than 1 every 3 seconds, with most of their comments (88%) consisting only of emojis. It’s clear from these initial statistics that there is automated activity on the platform, but these methods can only produce a rough lower bound on the extent of automated activity on the platform. We therefore hope future work can narrow our bounds on the amount of inauthentic activity on the platform and in this dataset.
3 Discussion
The data we report in this work is critical for calibrating our understanding of work on TikTok. For example, when we look at the invasion of Ukraine on TikTok [30], understanding that no TikTok posts are tagged with a Russian origin, and that posts from Ukraine have an average view count (4.1k) of approximately half way between the global average (2.2k) and that of the USA (6.3k) is useful context. When looking at the presence of the global south on TikTok [22], learning that posts from countries in the global south [31] account for 81% of the video posts, while only obtaining an average view count of 1.5k adds scale to our understanding.
For work on Palestinian activism on TikTok [32], we can show that 0.02% of the videos in our hour dataset were classified as a Palestine protest associated topic, but those videos had a significantly lower (p < 0.02) average view count of 440 views compared to the hour dataset view average of 2.5k. For work looking at the role of war on TikTok[33], we can show that 0.15% of the videos in our hour dataset were classified as a military/soldier associated topic, and those videos received an average of 5.2k views. And for work looking at the link between TikTok and public health [34], we can add context by showing that videos associated with the topic ‘Medical Examination and Healthcare Setting’ had a significantly higher (p < 0.02) average view mean of 28.8k views than the average.
In order to allow others to use the data we have collected here to calibrate their studies, we have released aggregate data distributions. We understand that we won’t cover every use case with this data release, so we invite interested parties to contact us for additional specific statistics.
Our dataset is not perfect. Collecting the data for this dataset took 5 months, meaning that the first batch of data was collected 2 months after the day chosen, and the last batch of data was collected 7 months after the day chosen. This introduces temporal bias into our dataset, where older videos will accumulate more engagement, and are also more likely to be taken down, or deleted. However, looking at our comment engagement offset (Fig. 6), we can see that the median time of engagement for all comment authors is around 29.9 hours after the video was posted. It seems likely that the median time from video post to comment engagement is similar to the median time offset for other engagement, meaning that at least for engagement, the temporal bias should have minimal effect. We can see from Figure 2(b) that there is a decrease in video count over the hour we collected down to 83.7% that of the beginning of the hour, which could be due to temporal collection bias, true post volume decrease, or other factors.
We accepted the slow collection rate of our method due to its ability to give us a full representative slice of TikTok. However, producing fully geographically unbiased statistics ideally requires a full 24 hour’s worth of data is collected [17]. This is why we collected our second dataset of a minute from each hour. But being able to see the full spectrum of activity would necessarily require a full 24 hours worth of data.
We also cannot speak to the proportion of authentic to inauthentic content in this dataset. Detecting bots and influence operations is a pressing issue on TikTok, but is still difficult. We know that bots exist on the platform [35], but it is unlikely that existing bot detection methods transfer well [13], due to the difference in text distribution and platform properties. In this dataset, we do not have full chronological data timelines to allow methods that exploit timestamps to work fully [36], or network data to allow graph methods to work [37]. While we used simple heuristics to examine the least sophisticated types of bots on the platform, in the age of LLMs and advanced influence operations, more comprehensive methods must be used to be able to detect this activity, which we will leave for future work.
Having said this, we believe the statistics we have presented here to be of high value to the community. With this work we have presented the first large vertical slice of posts from TikTok, and delivered the essential statistics to help understand the platform. Our data release will allow those doing studies of TikTok to contextualize their findings, and work to build on understanding of this global platform.
4 Methods
4.1 ID Generation
Every entity (video, user, and comment) on TikTok has a unique ID attached, which takes the form of a very large integer. In order to create an ID generator, we need to know how these integers are created for every new video.
We have a starting point, which is that we know that if we convert the integer to a 64 bit binary representation, we can take the first 32 bits, convert these 32 bits back to an integer representation, we get a unix timestamp, which tends to be within seconds of the recorded creation time of the video recorded in the video metadata [38]. This indicates that within the total 64 bits of the ID, there is encoded information. Many large scale distributed systems that need to generate globally unique IDs, use systems similar to the Snowflake system [39, 40, 41, 42]. If the ID generation method is similar to snowflake, then we would expect that the rest of the bits are some combination of milliseconds data, and data relating to TikTok’s infrastructure (worker IDs, sharding IDs, sequence IDs etc.). If we can understand what composes the rest of these bits, when generating IDs, we can reduce the search space from $4.3\times 10^{9}$ combinations down to a manageable number.
Domingues et al. [43] found that the $3^{rd}$ lowest hexadecimal value determines the type of the ID (post, user, comment etc.), meaning the $52^{nd}$ to the $56^{th}$ bits [43].
We scraped a TikTok video metadata dataset of 225 thousand videos, attempting to capture videos that were likely produced in a variety of places to get high geographical distribution. The first method we used for this was datasets associated with accounts known to be in a specific country, which were available for Canada [44] and Germany [45]. For countries where we were unable to find country specific seedlists, we collected videos using the hashtag for the name of the country in the country’s official language. As long as a reasonable percentage of videos with the hashtag do indeed originate in that country, by looking at ID distribution, we can estimate whether a good ID coverage for that geographical region.
Country Method No. Videos Estimated Coverage
Canada Seedlist 118994 99.93%
Germany Seedlist 35256 99.94%
Nigeria Country Name Search 41127 99.86%
Indonesia Country Name Search 19509 99.83%
Brazil Country Name Search 14079 99.77%
All 225127 99.97%
Table 3: We estimate the percentage of videos we can capture, given the unique ID patterns we have found, using the Good-Turing estimator [46].
We then take the numerical IDs for those videos, and convert them to 64 bits of binary. We confirm that the first 32 bits of the ID correspond to the creation time of the video. This further implies that the TikTok ID generation system uses a Snowflake like method. We know that it isn’t exactly using the snowflake method, as it uses the first bit as part of the timestamp (Snowflake reserves this first bit as always zero) [39].
73414563485943104010110010111100010000101111010011100000000000000000000110100000001Convert to binary1709316007Convert to base 1019:00:07UTC 1/3/2024MillisecondsSequence IDsEntity TypeMachine IDs
Figure 10: Inferred ID composition
To gather more information about the rest of the bits in the IDs, we found the inter-bit correlations by training logistic regression models to predict each of the bits in the IDs, given the rest of the bits as inputs. We found that bits 32 to 41 had no relation to each other, but that bits 42 to 63 had strong correlations, implying that there are a smaller set of unique values for the later part of this ID. Additionally, we see two clear sections in this later part, implying two different mechanisms for setting them, potentially worker IDs or sharding IDs.
We created a histogram of the first section, and found that values between 0 and 999 are roughly equally likely, with only 0.0004% of values between 1000 to 1023, as shown in Figure 11(a). This implies that this value is the millisecond section of the ID. The few other values are likely from other ID formatting mechanisms, or non video-type posts such as livestreams.
(a) Value occurrence distribution for bits 32 to 42.
(b) Distribution of occurrence of values in bits 58 to 64. The different in distributions between geographic areas implies that this section of the ID encodes information related to the datacentre in which the ID was created, matching with the machine ID in Snowflake IDs [39].
(c) Frequency of value occurrence in bits 45 to 50. There is minimal difference in the distributions between countries, implying this section of the ID is zone agnostic.
(d) Visualization of the potential counter section of the ID, for a second of collected data, and two machine IDs.
Figure 11: Representations of frequency of bits in the video IDs.
We also made histograms of the later section of the ID, and found that IDs were highly concentrated into specific values.
The original Snowflake ID system contains a machine specific ID in the generated ID, where there are a finite number of machines that generate new IDs. In a global distributed system, it’s typically desired that different geographic regions (i.e. continents) are serviced by regional datacentres in order to reduce latency [47], and therefore machine IDs will likely be region specific. For that reason, we collected datasets correlated with globally distributed countries. We used Canada and Germany to cover North America and Europe as we had pre-existing, large datasets of videos from those countries. We searched for large country specific pre-existing datasets for other countries, but to the best of our knowledge finding an absence of such datasets, collected datasets by getting videos using the hashtag of the name of those countries, with the intuition that most videos featuring those hashtags will have been created in those countries. Due to the hashtag search mechanism being limited to 1000 videos, we then collected videos from the accounts discovered via the hashtag method to obtain more videos.
To ensure we have a representative dataset, we wanted to create some statistical assurances that we have captured enough of the possible ID patterns that we can say with confidence we are capturing the vast majority of videos. To do this, we can turn to the Good-Turing estimator [46], which, given our existing dataset of IDs, where we treat the non time encoding bits as the type of the ID, we can then estimate the number of each type of ID we have to estimate the probability of finding a previously undiscovered type of ID. This probability is then the probability of a video existing that we are not able to capture given our current technique. We report the corresponding percentage of coverage in Table 3.
We see differences in the IDs in the later section as shown in Figure 11(b). The difference in distributions between geographic areas implies that this section of the ID encodes information related to the datacentre in which the ID was created, which corresponds to the machine ID in Snowflake IDs [39].
Finally we wanted to see if we could find sequence IDs, or counters, in the ID, as they are present in other distributed ID systems [39]. The only section that we have not yet established a reasonable inference of function for is bits 42 to 52. When comparing different ID types (video post, user, comment), we see differences in the distribution in these bits. However, there may also be a counter representation in this section. If found, and if they increment predictably, this would allow us to reduce the number of potential IDs that we need to try when doing ID generation. For video IDs, we see minimal variation in bits 42 to 52 other than in bits 45 to 50, for which we plot the distribution in Figure 11(c).
We did find evidence of counters in the IDs, but they do not count up uniformly, sometimes missing steps, as shown in Figure 11(d). This means we cannot use them to narrow down our ID search space.
We then needed to generate these IDs, and request them from TikTok to see if they correspond to valid videos. To do this, we simply find all the possible variations of bit sequences for bits 42 to 63 (all the bits past the millisecond segment), and request each sequence of these for every millisecond. This results in 504 IDs to check for each millisecond.
Geographical Coverage
The Good-Turing estimator gives us confidence in what percentage of videos we have for each geographical area, but the question remains if we have good geographic coverage. After looking at the geographic distribution of the videos collected, we saw that the notable places where video counts were lower than others were East Asia, Oceania, India, and Russia. Though there are good reasons to believe that video counts in India and Russia would naturally be lower due to the ban of TikTok in those countries [23, 48], we collected extra datasets correlated with those regions (India, Australia, Japan, Russia) to check if there were previously undiscovered IDs, see details in Table 4. We found that the unique IDs we had previously collected covered at least 90% of the videos from those areas, and for India and Japan, at least 95%, which we feel is sufficient to ensure a dataset is representative of those areas.
Country Hashtag Name No. Videos Estimated Coverage Fetch Overlap
Russia Россия 10288 99.35% 91%
Japan {CJK}UTF8min日本 9363 99.30% 96%
Australia australia 6522 99.13% 92%
India india 6649 99.25% 97%
Table 4: We estimate the percentage of videos we can capture, given the unique ID patterns we have found, using the Good-Turing estimator [46].
Specifically we notice that there are no video posts attributed to Russia in our dataset, which is notable given that both India and Russia have banned TikTok, but we still obtain posts attributed to India. To ensure that we have not completed missed ID aspects that are associated with Russia, we collected the full metadata of the 9% of posts we collected under the Russian hashtag that would not have been covered by our original fetch. We saw that none of them were attributed to Russia, only a range of other countries, implying that Russians using TikTok, or those ostensibly posting from Russia, are using VPNs, a common practice in other areas with banned internet services [49].
4.1.1 Future Work
We investigated the apparent counters found in the IDs, and trialled methods to predict the next valid video ID given previous counters. However, in the end we decided we wanted to get as complete a sample as we could, and therefore just requested all possible ID combinations. But if a method could be successfully developed here, it would greatly cut down the number of video IDs that would need to be requested, expanding the number of videos that could be collected through this method.
If we accepted a more probabilistic approach to video collection, we could have collected many more views for many fewer requests. However, this would have meant a less complete slice of TikTok. If from the data we have gathered, we can determine that ID patterns with more frequent video hits are not meaningfully different than ID patterns with less frequent occurrence of video posts, we could confidently roll out such a system.
We also looked at comment and user IDs, and a similar method used here could be applied to gather a sample of users. A random sample of comments would be more difficult due to the query for them requiring the video ID as well as the comment ID.
4.2 Collection
Once we could generate a full set of IDs, we needed to request them from TikTok. The typical way of viewing a TikTok video is by going to the URL: https://tiktok.com/@{username}/video/{video_id}. But we discovered that one can use an arbitrary username in the URL, and if the video corresponding to the video ID exists, one will be re-directed to the existing video page.
In total, 0.8% of the IDs we requested returned either an available video, or an error indicating that the video had at some point existed.
We collected two datasets, one of all video posts from 5-6pm UTC on the 10th April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all videos from the $42^{nd}$ minute of each hour from 10pm UTC on the 9th April, till 10pm on the 10th April, to obtain across day representative data. The entire collection process took 5 months.
For the actual video collection process, we use Dask Distributed[50], running on a cluster of Raspberry Pis managed by SLURM [51].
4.3 Global Usage Estimation
We correct post count values for each country by using our country coverage estimates from section 4.1 to obtain regional coverage estimates, and applying regional coverage estimates to each country in that region (for example, correcting post counts for New Zealand and other Pacific countries by assuming the same coverage as for Australia).
4.4 Videos
4.4.1 Topics
We embedded each video clip in a latent space using the xclip-base-patch32 model  [^8] https://huggingface.co/microsoft/xclip-base-patch32 [52].
We then use the BERTopic  [^9] https://maartengr.github.io/BERTopic/index.html [53] library to cluster the video embeddings, then use the InternVL2_5-2B  [^10] https://huggingface.co/OpenGVLab/InternVL2_5-2B [54] vision-language model to obtain captions for representative documents from each cluster, and finally Phi-4-mini-instruct  [^11] https://huggingface.co/microsoft/Phi-4-mini-instruct [55] to summarize clusters. We then use datamapplot  [^12] https://datamapplot.readthedocs.io/en/latest/ to plot the resulting clusters and descriptions.
4.4.2 Classification of videos featuring children
We needed to build a classifier to classify each video in our 24 hour dataset as containing a child or not. The xclip video embeddings were used as the representations of each video. To train and test a classifier, we used a dataset of TikTok accounts [56]. The dataset contains 5,248 unique videos depicting children, including babies and young teenagers below the age of 13, and an additional 6,972 videos classified as not containing children. Videos were collected through a snowball sampling approach based on child-related keywords and are not restricted to a certain language or region. We evaluated methods via F1 on a 5-fold cross-validation train/test split. We tested several classifier methods, and found a stacking ensemble method using a logistic regression, histogram-based gradient boosting classifier  [^13] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.HistGradientBoostingClassifier.html, and an extra trees classifier [57] produced the best mean F1 of 0.86.
Given the difficulty of detecting children of varying age groups in contexts, the obtained F1 score of 0.86 seems particularly dependable. As stated by Schirmer et al. [56], the range of contexts in which videos feature children is vast: Videos contain vlogs of new mothers showing their newborn, parents playing with their toddlers, and families performing popular TikTok dances. Further, children might be shown only for a couple of seconds or throughout the whole video. Therefore, textual descriptions of videos can contain a lot of noise or might even miss child-related cues, making the performance of our classifier even more promising.
5 Data Availability
We have released aggregated statistics from our collected dataset in order to preserve the privacy of those in our dataset. We have released the data here URL.
6 Code Availability
The code for this work is available at URL
[…]
Supplementary information
If your article has accompanying supplementary file/s please state so here.
Authors reporting data from electrophoretic gels and blots should supply the full unprocessed scans for key as part of their Supplementary information. This may be requested by the editorial team/s if it is missing.
Please refer to Journal-level guidance for any specific requirements.